OpenStudy (anonymous):

A post is 9 ft from one wall of a room and 40 ft from the wall at right angle to it. How many ft is it from the intersection of the two walls?

sam (.sam.):

Let one wall be the x-axis and let the other wall be the y-axis. The post is at \((4,9)\), the intersection of the walls is \((0,0)\) ---------------------------------------------------------- Distance from (0,0) to (4,9): \[d = \sqrt{4^2+9^2}= \sqrt{16+36}= \sqrt{52} ft\]

OpenStudy (anonymous):

Choices: a. 7 b. 21 c. 41 d. 49

sam (.sam.):

Can you use a calculator?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Aah, okay. Thank You :D

sam (.sam.):

Wait

sam (.sam.):

The post is at (40,9), the intersection of the walls is (0,0) ------------------------------------------------ Distance from (0,0) to (40,9): \[d = \sqrt{40^2+9^2}= \sqrt{1600+81}= 41 ft\]

sam (.sam.):



OpenStudy (anonymous):

Yeah, that's what i mean.

sam (.sam.):

Okay, that's done